OpenStudy (anonymous):

If f(x) = ax+b where a and b are constants, find a and b for f(1) = 7 and f(3) = -5. Please help

OpenStudy (ybarrap):

f(1) = 7 = a(1) + b f(3) = -5 = a(3) + b f(1) - f(3) => 7 - (-5) = 12 = a(1) - a(3) = a(1 - 3) = a(-2) So a(-2) = 12 a = 12/(-2) a = -6 using f(1) to find b b = 7 - a(1) = 7 - (-6)(1) = 7 + 6 = 13 Ans. f(x) = -6x + 13

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (ybarrap):

np